# On the Trade-off between Flatness and Optimization in Distributed Learning

### v1

@@ Acknowledgments
Substituting (F) into (386), we obtain
@@
Substituting (410), (412) and (414) into (F), we then have
@@
Substituting (F) into (416), we obtain
@@
Substituting (450) and (426) into (F), we obtain
@@
where $(a)$ follows from (36). We focus on the first term in (F) as it is on the order of $O(\frac{1}{\mu})$, which dominates the second term $O(1)$. According to (424), it holds that
@@
Substituting (459) into (F), we have
@@
where $(a)$ follows from (424) and (425), $(b)$ follows from (F).
Substituting (F) and (F) into (422), we obtain
@@
We further substitute (F) and (462) into (339), and obtain
$\mathds{E}\|{\widetilde{\boldsymbol{\scriptstyle\mathcal{W}}}}_{n}^{\prime}\|^{2}_{I\otimes\bar{H}}=$ $\frac{\mu K}{2B}\mathrm{Tr}\left(\left(I-(I-\mu\bar{H})^{2(n+1)}\right)\bar{R}_{s}\right)+\mu^{2}\|d^{\sf T}\mathcal{A}_{2}\mathcal{V}_{\alpha}(I-\mathcal{P}_{\alpha})^{-1}(I-\mathcal{P}_{\alpha}^{n+1})\|^{2}_{I\otimes\bar{H}}\pm O(\frac{\mu^{\frac{\gamma^{\prime}}{2}+1}}{B})\pm o(\mu^{1+\eta})+o(\mu^{2})$ (463)
@@
Basically, for the centralized method, we substitute (34) and (464) into (333), and obtain
$\mathrm{ER}_{n,cen}$ $=\frac{\mu}{4BK}\mathrm{Tr}\left(\left(I-(I-\mu\bar{H})^{2(n+1)}\right)\bar{H}\right)\pm O(\mu^{1.5\gamma^{\prime}})\pm O(\frac{\mu^{\frac{\gamma^{\prime}}{2}+1}}{B})+o(\mu^{1+\eta})$
=$\overset{(a)}{=}\frac{\mu}{4BK}\mathrm{Tr}\left(\left(I-(I-\mu\bar{H})^{2(n+1)}\right)\bar{H}\right)\pm O\left(\mu^{1.5(1+\eta)}\right)+o(\mu^{1+\eta})$
$=\frac{\mu}{4BK}\mathrm{Tr}\left(\left(I-(I-\mu\bar{H})^{2(n+1)}\right)\bar{H}\right)+o(\mu^{1+\eta})$ (465)
@@
For consensus, we apply $A_{2}=I$, and substitute (463) and (34) into (333), which gives
$\mathrm{ER}_{n,con}=$ $\frac{\mu}{4BK}\mathrm{Tr}\left(\left(I-(I-\mu\bar{H})^{2(n+1)}\right)\bar{H}\right)+\frac{\mu^{2}}{2K}\|d^{\sf T}\mathcal{V}_{\alpha}(I-\mathcal{P}_{\alpha})^{-1}(I-\mathcal{P}_{\alpha}^{n+1})\|^{2}_{I\otimes\bar{H}}\pm o(\mu^{2})\pm o(\mu^{1+\eta})$ (467)
@@
$\mathrm{ER}_{n,dif}=$ $\frac{\mu}{4BK}\mathrm{Tr}\left(\left(I-(I-\mu\bar{H})^{2(n+1)}\right)\bar{H}\right)+\frac{\mu^{2}}{2K}\|d^{\sf T}\mathcal{V}_{\alpha}\mathcal{P}_{\alpha}(I-\mathcal{P}_{\alpha})^{-1}(I-\mathcal{P}_{\alpha}^{n+1})\|^{2}_{I\otimes\bar{H}}\pm o(\mu^{2})\pm o(\mu^{1+\eta})$ (469)
Finally, we specify the number of iterations required for the algorithms to escape the local basin. For the centralized method, recall (F), and let $\sigma_{i}$ with $i=1,\ldots,M$ denote the i-th largest eigenvalue of $\bar{H}$, we have
$\mathrm{ER}_{n,cen}$ $=\frac{\mu}{4BK}\mathrm{Tr}\left(\left(I-(I-\mu\bar{H})^{2(n+1)}\right)\bar{H}\right)$
=$\overset{(a)}{=}\frac{\mu}{4BK}\mathrm{Tr}\left(\left(I-e^{-2\mu(n+1)\bar{H}}\right)\bar{H}\right)\pm o\left(\frac{\mu}{B}\right)$
$=\frac{\mu}{4BK}\sum\limits_{i=1}^{M}\sigma_{i}(1-e^{-2\mu(n+1)\sigma_{i}})\pm o\left(\frac{\mu}{B}\right)$
=$\overset{(b)}{=}\frac{\mu c}{4BK}\sigma_{1}(1-e^{-2\mu(n+1)\sigma_{1}})\pm o\left(\frac{\mu}{B}\right)$ (470)
@@
and (b) follows from the fact that for any positive-definite matrix $X$, we have
$\mathrm{Tr}(X)=c\lambda_{\max}(X)$ (473)
with $1\leq c\leq M$, where $\lambda_{\max}$ is the largest eigenvalue of $X$. Let $\mathrm{ER}_{n,cen}=h$, then we need to solve the following equation associated with $n$:
$\frac{\mu c}{4BK}\sigma_{1}(1-e^{-2\mu(n+1)\sigma_{1}})=h\pm o\left(\frac{\mu}{B}\right)$ (474)
@@
$n=\frac{1}{2\mu\sigma_{1}}\mathrm{log}\frac{1}{1-\frac{4BK(h\pm o\left(\frac{\mu}{B}\right))}{\mu c\sigma_{1}}}-1=O(\frac{1}{\mu})$ (475)
@@
Appendix G Proof for Corollary III.9: Comparison between consensus and diffusion
@@
Appendix H Proof for Corollary IV.1
In this section, we examine the closed-form excess-risk performance of the centralized, consensus and diffusion methods in the long term when $n\to\infty$ which corresponds to their optimization performance. To achieve the goal, we first need to verify how accurately the approximate model defined in (48) can approximate the original model in (42). To do this, similar to Lemmas III.6 and III.7, we need to analyze the upper bounds of $\mathds{E}\|{\widetilde{\boldsymbol{\scriptstyle\mathcal{W}}}}_{n}\|^{2}$, $\mathds{E}\|{\widetilde{\boldsymbol{\scriptstyle\mathcal{W}}}}_{n}\|^{2}$ and $|\mathds{E}\|\widetilde{\boldsymbol{\scriptstyle\mathcal{W}}}_{n}^{\prime}\|^{2}-\mathds{E}\|\widetilde{\boldsymbol{\scriptstyle\mathcal{W}}}_{n}\|^{2}|$, for which the following two lemmas can be established:
Lemma H.1.
@@
Lemma H.2.
(Deviation bounds of the centralized method in the long term). Under the same conditions of Lemma H.1, after sufficient iterations, the second and fourth-order moments of ${\widetilde{\boldsymbol{\scriptstyle\mathcal{W}}}}_{n}$, the second-order moment of ${\widetilde{\boldsymbol{\scriptstyle\mathcal{W}}}}_{n}^{\prime}$, and the approximation error of the short term model, related to the centralized method are guaranteed to be upper bounded. Basically, it holds that
@@
The Proofs for Lemmas H.1 and H.2 can be found in Appendix I.
Recall (333), Lemmas H.1 and H.2 guarantee that the excess-risk value $\mathrm{ER}_{n}$ satisfies the following equality:
@@
Similar to Appendix F , we examine the size of $\mathds{E}\|{\widetilde{\boldsymbol{\scriptstyle\mathcal{W}}}}_{n}^{\prime}\|^{2}_{I\otimes\bar{H}}$, where (339) is still valid. First, we know from Theorem 9.3 in [6] that $\mathcal{C}$ is a stable matrix, thus when $n\to\infty$, we have:
@@
By following the same derivation process of (340)–(F), (F) is sill valid. Specifically, since $P_{\alpha}$ is a diagonal matrix where the absolute values of all elements are smaller than 1, in the steady state when $n\to\infty$, we have
@@
We finally derive closed-form expressions for the excess risk in the long term when $n\to\infty$. Basically, for the centralized method, we substitute (34), (339), (519), (520) and $\xi=1+\eta$ into (516), and obtain
$\mathrm{ER}_{\infty,cen}$ $=\frac{\mu}{4BK}\mathrm{Tr}\left(\bar{H}\right)\pm O\left(\mu^{1.5(1+\eta)}\right)$ (521)
As for decentralized methods, we separately show the results of consensus and diffusion. That is, for consensus, we substitute (34), (339), (518), (520), $\xi=\gamma$ and $A_{2}=I$ into (516), and obtain
$\mathrm{ER}_{\infty,con}=$ $\frac{\mu}{4BK}\mathrm{Tr}\left(\bar{H}\right)+\frac{\mu^{2}}{2K}\|d^{\sf T}\mathcal{V}_{\alpha}(I-\mathcal{P}_{\alpha})^{-1})\|^{2}_{I\otimes\bar{H}}\pm O(\mu^{1.5\gamma})\pm o(\mu^{2})$ (522)
Similarly, for diffusion, we substitute (34), (339), (518), (520), $\xi=\gamma$ and $A_{2}=A$ into (516), and obtain
$\mathrm{ER}_{\infty,dif}=$ $\frac{\mu}{4BK}\mathrm{Tr}\left(\bar{H}\right)+\frac{\mu^{2}}{2K}\|d^{\sf T}\mathcal{V}_{\alpha}\mathcal{P}_{\alpha}(I-\mathcal{P}_{\alpha})^{-1})\|^{2}_{I\otimes\bar{H}}\pm O(\mu^{1.5\gamma})\pm o(\mu^{2})$ (523)
@@
Appendix I Proof for Lemmas H.1 and H.2
The proofs for Lemmas H.1 and H.2 are similar to Appendices CE, and also similar to the proofs in [6].
I-A Proof for $\mathds{E}\|{\widetilde{\boldsymbol{\scriptstyle\mathcal{W}}}}_{n}\|^{2}$.
Assume $J(w)$ is locally $\nu$-strongly-convex around $w^{\star}$. We first verify the bounds for $\mathds{E}\|{\widetilde{\boldsymbol{\scriptstyle\mathcal{W}}}}_{n}\|^{2}$: By following the same logic of (104)–(C), we substitute the strongly-convex condition into (C), and choose $10<t=1-\mu\nu<$ 1, then we obtain
$\mathds{E}\|\mathcal{V}^{\sf T}{\widetilde{\boldsymbol{\scriptstyle\mathcal{W}}}}_{n}\|^{2}=\mathds{E}\left[\begin{array}[]{c}\|\bar{\boldsymbol{w}}_{n}\|^{2}\\ \|\check{\boldsymbol{w}}_{n}\|^{2}\end{array}\right]\leq\left[\begin{array}[]{cc}1-\mu\nu+O(\mu^{2})&O(\mu)\\ O(\mu^{2})&\rho(P_{\alpha})+O(\mu^{2})\end{array}\right]\left[\begin{array}[]{c}\mathds{E}\|\bar{\boldsymbol{w}}_{n-1}\|^{2}\\ \mathds{E}\|\check{\boldsymbol{w}}_{n-1}\|^{2}\end{array}\right]+\left[\begin{array}[]{c}O(\frac{\mu^{2}}{B})\\ O(\mu^{2})\end{array}\right]$ (535)
@@
$\lim_{n\to\infty}\mathds{E}\|\mathcal{V}^{\sf T}{\widetilde{\boldsymbol{\scriptstyle\mathcal{W}}}}_{n}\|^{2}$ $\leq\Gamma_{3}^{n+1}\mathds{E}\|\mathcal{V}^{\sf T}{\widetilde{\boldsymbol{\scriptstyle\mathcal{W}}}}_{-1}\|^{2}+\sum\limits_{i=0}^{n}\Gamma_{3}^{i}\left[\begin{array}[]{c}O(\frac{\mu^{2}}{B})\\ O(\mu^{2})\end{array}\right]=(I-\Gamma_{3})^{-1}\left[\begin{array}[]{c}O(\frac{\mu^{2}}{B})\\ O(\mu^{2})\end{array}\right]$ (545)
@@
$\mathds{E}\|{\widetilde{\boldsymbol{\scriptstyle\mathcal{W}}}}_{n}\|^{2}\leq\|\mathcal{V}\|^{2}\mathds{E}\|{\mathcal{V}^{\sf T}\widetilde{\boldsymbol{\scriptstyle\mathcal{W}}}}_{n}\|^{2}\leq O(\frac{\mu}{B})+O(\mu^{2})=O(\mu^{\gamma})$ (553)
@@
I-B Proof for $\mathds{E}\|{\widetilde{\boldsymbol{\scriptstyle\mathcal{W}}}}_{n}\|^{4}$.
The derivation process is similar to Appendix D. The only difference is in (D) for the strongly-convex case, if we choose $t=1-\mu\nu$, it holds that
@@
Similarly, for the centralized method, we substitute (554) into (I-B) and iterate it, and obtain
@@
[…]
I-C Proof for $|\mathds{E}\|\widetilde{\boldsymbol{\scriptstyle\mathcal{W}}}_{n}^{\prime}\|^{2}-\mathds{E}\|\widetilde{\boldsymbol{\scriptstyle\mathcal{W}}}_{n}\|^{2}|$ and $\mathds{E}\|{\widetilde{\boldsymbol{\scriptstyle\mathcal{W}}}}_{n}^{\prime}\|^{2}$.
We finally analyze the bounds of $|\mathds{E}\|\widetilde{\boldsymbol{\scriptstyle\mathcal{W}}}_{n}^{\prime}\|^{2}-\mathds{E}\|\widetilde{\boldsymbol{\scriptstyle\mathcal{W}}}_{n}\|^{2}|$ and $\mathds{E}\|{\widetilde{\boldsymbol{\scriptstyle\mathcal{W}}}}_{n}^{\prime}\|^{2}$. The proofs are similar to Appendix E. The only difference is that we now consider the long-term properties of $\mathds{E}\|\widetilde{\boldsymbol{\scriptstyle\mathcal{W}}}_{n}^{\prime}\|^{2}-\mathds{E}\|\widetilde{\boldsymbol{\scriptstyle\mathcal{W}}}_{n}\|^{2}|$ when $n\to\infty$. Recalling (298) and (301), after enough iterations when $n$ is sufficiently large, we have
$\mathds{E}\|\mathcal{V}_{\alpha}^{\sf T}{{\boldsymbol{{z}}}}_{n}\|^{2}=(I-\Gamma^{\prime})^{-1}\left[\begin{array}[]{c}O(\mu)\mathds{E}\|\widetilde{\boldsymbol{\scriptstyle\mathcal{W}}}_{n-1}\|^{4}\\ O(\mu^{2})\mathds{E}\|\widetilde{\boldsymbol{\scriptstyle\mathcal{W}}}_{n-1}\|^{4}\end{array}\right]$ (588)
Substituting (306) and (583) into (588), we have
$\mathds{E}\left[\begin{array}[]{c}\|\check{\boldsymbol{z}}_{n}\|^{2}\\ \|\bar{\boldsymbol{z}}_{n}\|^{2}\end{array}\right]\leq$ $\left[\begin{array}[]{cc}O(\frac{1}{\mu})&O(1)\\ O(\mu)&O(1)\end{array}\right]\left[\begin{array}[]{c}O(\mu^{2\gamma+1})\\ O(\mu^{2\gamma+2})\end{array}\right]=O(\mu^{2\gamma})$ (595)

### v2

@@ Acknowledgments
Note that according to (407) and (E), we have
$\left\|\sum\limits_{i=0}^{n}\mathcal{C}^{i}\mathcal{A}_{2}d\right\|=O(1)$ (431)
Then we have
$\left|\mu\mathds{E}\left(\sum\limits_{i=0}^{n}\mathcal{C}^{i}\mathcal{A}_{2}d\right)^{\sf T}(I\otimes\bar{H})(\mathcal{C}^{n+1}{\widetilde{\boldsymbol{\scriptstyle\mathcal{W}}}}_{-1}^{\prime})+\mathds{E}\|\mathcal{C}^{n+1}{\widetilde{\boldsymbol{\scriptstyle\mathcal{W}}}}_{-1}^{\prime}\|_{I\otimes\bar{H}}^{2}\right|$
$\leq\mu\mathds{E}\left\|\sum\limits_{i=0}^{n}\mathcal{C}^{i}\mathcal{A}_{2}d\right\|\left\|I\otimes\bar{H}\right\|\left\|\mathcal{C}^{n+1}\right\|\|{\widetilde{\boldsymbol{\scriptstyle\mathcal{W}}}}_{-1}^{\prime}\|+\|\mathcal{C}^{n+1}\|^{2}\|\bar{H}\|\mathds{E}\|{\widetilde{\boldsymbol{\scriptstyle\mathcal{W}}}}_{-1}^{\prime}\|^{2}$
$\overset{(a)}{\leq}O(\mu)\times o(\sqrt{\frac{\mu}{B}})+o(\frac{\mu}{B})$
=$\overset{(b)}{=}o(\mu^{0.5(3+\eta)})+o(\mu^{1+\eta})$ (432)
where $(a)$ follows from Assumption III.1, (408) and (431), and $(b)$ follows from (48), then we have:
Substituting (E) into (433), we obtain
@@
Substituting (438), (440) and (442) into (E), we then have
@@
Substituting (E) into (444), we obtain
@@
Substituting (478) and (454) into (E), we obtain
@@
where $(a)$ follows from (35). We focus on the first term in (E) as it is on the order of $O(\frac{1}{\mu})$, which dominates the second term $O(1)$. According to (452), it holds that
@@
Substituting (487) into (E), we have
@@
where $(a)$ follows from (452) and (453), $(b)$ follows from (E).
Substituting (E) and (E) into (450), we obtain
@@
We further substitute (E) and (490) into (364), and obtain
$\mathds{E}\|{\widetilde{\boldsymbol{\scriptstyle\mathcal{W}}}}_{n}^{\prime}\|^{2}_{I\otimes\bar{H}}=$ $\frac{\mu K}{2B}\mathrm{Tr}\left(\left(I-(I-\mu\bar{H})^{2(n+1)}\right)\bar{R}_{s}\right)+\mu^{2}\|d^{\sf T}\mathcal{A}_{2}\mathcal{V}_{\alpha}(I-\mathcal{P}_{\alpha})^{-1}(I-\mathcal{P}_{\alpha}^{n+1})\|^{2}_{I\otimes\bar{H}}$
±$\pm O(\frac{\mu^{\frac{\gamma^{\prime}}{2}+1}}{B})\pm o(\mu^{1+\eta})+o(\mu^{2})$ (491)
@@
Basically, for the centralized method, we substitute (492) into (359), and obtain
$\mathrm{ER}_{n,cen}$ $=\frac{\mu}{4B}\mathrm{Tr}\left(\left(I-(I-\mu\bar{H})^{2(n+1)}\right)\bar{R}_{s}\right)\pm O(\mu^{1.5\gamma^{\prime}})\pm O(\frac{\mu^{\frac{\gamma^{\prime}}{2}+1}}{B})+o(\mu^{1+\eta})$
=$\overset{(a)}{=}\frac{\mu}{4B}\mathrm{Tr}\left(\left(I-(I-\mu\bar{H})^{2(n+1)}\right)\bar{R}_{s}\right)\pm O\left(\mu^{1.5(1+\eta)}\right)+o(\mu^{1+\eta})$
$=\frac{\mu}{4B}\mathrm{Tr}\left(\left(I-(I-\mu\bar{H})^{2(n+1)}\right)\bar{R}_{s}\right)+o(\mu^{1+\eta})$ (493)
@@
For consensus, we apply $A_{2}=I$, and substitute (E) into (359), which gives
$\mathrm{ER}_{n,con}=$ $\frac{\mu}{4B}\mathrm{Tr}\left(\left(I-(I-\mu\bar{H})^{2(n+1)}\right)\bar{R}_{s}\right)+\frac{\mu^{2}}{2K}\|d^{\sf T}\mathcal{V}_{\alpha}(I-\mathcal{P}_{\alpha})^{-1}(I-\mathcal{P}_{\alpha}^{n+1})\|^{2}_{I\otimes\bar{H}}\pm o(\mu^{2})\pm o(\mu^{1+\eta})$ (495)
@@
$\mathrm{ER}_{n,dif}=$ $\frac{\mu}{4B}\mathrm{Tr}\left(\left(I-(I-\mu\bar{H})^{2(n+1)}\right)\bar{R}_{s}\right)+\frac{\mu^{2}}{2K}\|d^{\sf T}\mathcal{V}_{\alpha}\mathcal{P}_{\alpha}(I-\mathcal{P}_{\alpha})^{-1}(I-\mathcal{P}_{\alpha}^{n+1})\|^{2}_{I\otimes\bar{H}}\pm o(\mu^{2})\pm o(\mu^{1+\eta})$ (497)
Recalling (358), the common $O(\frac{\mu}{B})$ term in (E), (495), and (497), associated with gradient noise, corresponds to the size of the term in (358) that is related to the network centroid. In addition, the additional $O(\mu^{2})$ term in the escaping efficiency of decentralized methods arises due to the consensus distance.
Finally, we specify the number of iterations required for the algorithms to escape the local basin. For the centralized method, recall (E), and let $\sigma_{i}$ with $i=1,\ldots,M$ denote the i-th largest eigenvalue of $\bar{H}$, we have
$\mathrm{ER}_{n,cen}$ $=\frac{\mu}{4B}\mathrm{Tr}\left(\left(I-(I-\mu\bar{H})^{2(n+1)}\right)\bar{R}_{s}\right)$
=$\overset{(a)}{=}\frac{\mu}{4B}\mathrm{Tr}\left(\left(I-e^{-2\mu(n+1)\bar{H}}\right)\bar{R}_{s}\right)\pm o\left(\frac{\mu}{B}\right)$ (498)
@@
Note that
$\mathrm{Tr}\left(I-e^{-2\mu(n+1)\bar{H}}\right)\lambda_{\min}(\bar{R}_{s})\leq\mathrm{Tr}\left(\left(I-e^{-2\mu(n+1)\bar{H}}\right)\bar{R}_{s}\right)\leq\mathrm{Tr}\left(I-e^{-2\mu(n+1)\bar{H}}\right)\lambda_{\max}(\bar{R}_{s})$ (501)
where $\lambda_{\min}(\bar{R}_{s})$ and $\lambda_{\max}(\bar{R}_{s})$ denote the smallest and largest eigenvalues of $\bar{R}_{s}$, respectively. Thus, we have:
$\mathrm{Tr}\left(\left(I-e^{-2\mu(n+1)\bar{H}}\right)\bar{R}_{s}\right)=c_{1}\mathrm{Tr}\left(I-e^{-2\mu(n+1)\bar{H}}\right)=c_{1}\sum\limits_{i=1}^{M}(1-e^{-2\mu(n+1)\sigma_{i}})$ (502)
where
$\lambda_{\min}(\bar{R}_{s})\leq c_{1}\leq\lambda_{\max}(\bar{R}_{s})$ (503)
Also, for any positive-definite matrix $X$, we have
$\mathrm{Tr}(X)=c_{2}\lambda_{\max}(X)$ (504)
with $1\leq c_{2}\leq M$, where $\lambda_{\max}(X)$ is the largest eigenvalue of $X$. Let $c\overset{\Delta}{=}c_{1}c_{2}$, and substitute (502) and (504) into (E), we obtain
$\mathrm{ER}_{n}{=}\frac{\mu c}{4B}\sigma_{1}(1-e^{-2\mu(n+1)\sigma_{1}})\pm o\left(\frac{\mu}{B}\right)$ (505)
Let $\mathrm{ER}_{n,cen}=h$, then we need to solve the following equation associated with $n$:
$\frac{\mu c}{4B}\sigma_{1}(1-e^{-2\mu(n+1)\sigma_{1}})=h\pm o\left(\frac{\mu}{B}\right)$ (506)
@@
$n=\frac{1}{2\mu\sigma_{1}}\mathrm{log}\frac{1}{1-\frac{4B(h\pm o\left(\frac{\mu}{B}\right))}{\mu c\sigma_{1}}}-1=O(\frac{1}{\mu})$ (507)
@@
Appendix F Proof for Corollary III.8: Comparison between consensus and diffusion
@@
Appendix G Proof for Corollary IV.1
In this section, we examine the closed-form excess-risk performance of the centralized, consensus and diffusion methods in the long term when $n\to\infty$ which corresponds to their optimization performance. To achieve the goal, we first need to verify how accurately the approximate model defined in (47) can approximate the original model in (41). To do this, similar to Lemmas III.5 and III.6, we need to analyze the upper bounds of $\mathds{E}\|{\widetilde{\boldsymbol{\scriptstyle\mathcal{W}}}}_{n}\|^{2}$, $\mathds{E}\|{\widetilde{\boldsymbol{\scriptstyle\mathcal{W}}}}_{n}\|^{2}$ and $|\mathds{E}\|\widetilde{\boldsymbol{\scriptstyle\mathcal{W}}}_{n}^{\prime}\|^{2}-\mathds{E}\|\widetilde{\boldsymbol{\scriptstyle\mathcal{W}}}_{n}\|^{2}|$, for which the following two lemmas can be established:
Lemma G.1.
@@
Lemma G.2.
(Deviation bounds of the centralized method in the long term). Under the same conditions of Lemma G.1, after sufficient iterations, the second and fourth-order moments of ${\widetilde{\boldsymbol{\scriptstyle\mathcal{W}}}}_{n}$, the second-order moment of ${\widetilde{\boldsymbol{\scriptstyle\mathcal{W}}}}_{n}^{\prime}$, and the approximation error of the short term model, related to the centralized method are guaranteed to be upper bounded. Basically, it holds that
@@
The Proofs for Lemmas G.1 and G.2 can be found in Appendix H.
Recall (359), Lemmas G.1 and G.2 guarantee that the excess-risk value $\mathrm{ER}_{n}$ satisfies the following equality:
@@
Similar to Appendix E , we examine the size of $\mathds{E}\|{\widetilde{\boldsymbol{\scriptstyle\mathcal{W}}}}_{n}^{\prime}\|^{2}_{I\otimes\bar{H}}$, where (364) is still valid. First, we know from Theorem 9.3 in [6] that $\mathcal{C}$ is a stable matrix, thus when $n\to\infty$, we have:
@@
By following the same derivation process of (365)–(E), (E) is sill valid. Specifically, since $P_{\alpha}$ is a diagonal matrix where the absolute values of all elements are smaller than 1, in the steady state when $n\to\infty$, we have
@@
We finally derive closed-form expressions for the excess risk in the long term when $n\to\infty$. Basically, for the centralized method, we substitute (364), (551), (552) and $\xi=1+\eta$ into (548), and obtain
$\mathrm{ER}_{\infty,cen}$ $=\frac{\mu}{4B}\mathrm{Tr}\left(\bar{R}_{s}\right)\pm O\left(\mu^{1.5(1+\eta)}\right)$ (553)
As for decentralized methods, we separately show the results of consensus and diffusion. That is, for consensus, we substitute (364), (550), (552), $\xi=\gamma$ and $A_{2}=I$ into (548), and obtain
$\mathrm{ER}_{\infty,con}=$ $\frac{\mu}{4B}\mathrm{Tr}\left(\bar{R}_{s}\right)+\frac{\mu^{2}}{2K}\|d^{\sf T}\mathcal{V}_{\alpha}(I-\mathcal{P}_{\alpha})^{-1})\|^{2}_{I\otimes\bar{H}}\pm O(\mu^{1.5\gamma})\pm o(\mu^{2})$ (554)
Similarly, for diffusion, we substitute (364), (550), (552), $\xi=\gamma$ and $A_{2}=A$ into (548), and obtain
$\mathrm{ER}_{\infty,dif}=$ $\frac{\mu}{4B}\mathrm{Tr}\left(\bar{R}_{s}\right)+\frac{\mu^{2}}{2K}\|d^{\sf T}\mathcal{V}_{\alpha}\mathcal{P}_{\alpha}(I-\mathcal{P}_{\alpha})^{-1})\|^{2}_{I\otimes\bar{H}}\pm O(\mu^{1.5\gamma})\pm o(\mu^{2})$ (555)
@@
Appendix H Proof for Lemmas G.1 and G.2
The proofs for Lemmas G.1 and G.2 are similar to Appendices BD, and also similar to the proofs in [6].
H-A Proof for $\mathds{E}\|{\widetilde{\boldsymbol{\scriptstyle\mathcal{W}}}}_{n}\|^{2}$.
Assume $J(w)$ is locally $\nu$-strongly-convex around $w^{\star}$. We first verify the bounds for $\mathds{E}\|{\widetilde{\boldsymbol{\scriptstyle\mathcal{W}}}}_{n}\|^{2}$: By following the same logic of (95)–(B), we substitute the strongly-convex condition into (B), and choose $10<t=1-\mu\nu<$ 1, then we obtain
$\mathds{E}\left[\begin{array}[]{c}\|\bar{\boldsymbol{w}}_{n}\|^{2}\\ \|\check{\boldsymbol{w}}_{n}\|^{2}\end{array}\right]\leq\left[\begin{array}[]{cc}1-\mu\nu+O(\mu^{2})&O(\mu)\\ O(\mu^{2})&\rho(P_{\alpha})+O(\mu^{2})\end{array}\right]\left[\begin{array}[]{c}\mathds{E}\|\bar{\boldsymbol{w}}_{n-1}\|^{2}\\ \mathds{E}\|\check{\boldsymbol{w}}_{n-1}\|^{2}\end{array}\right]+\left[\begin{array}[]{c}O(\frac{\mu^{2}}{B})\\ O(\mu^{2})\end{array}\right]$ (567)
@@
$\lim_{n\to\infty}\mathds{E}\left[\begin{array}[]{c}\|\bar{\boldsymbol{w}}_{n}\|^{2}\\ \|\check{\boldsymbol{w}}_{n}\|^{2}\end{array}\right]$ $\leq\Gamma_{3}^{n+1}\mathds{E}\left[\begin{array}[]{c}\|\bar{\boldsymbol{w}}_{-1}\|^{2}\\ \|\check{\boldsymbol{w}}_{-1}\|^{2}\end{array}\right]+\sum\limits_{i=0}^{n}\Gamma_{3}^{i}\left[\begin{array}[]{c}O(\frac{\mu^{2}}{B})\\ O(\mu^{2})\end{array}\right]=(I-\Gamma_{3})^{-1}\left[\begin{array}[]{c}O(\frac{\mu^{2}}{B})\\ O(\mu^{2})\end{array}\right]$ (581)
@@
$\lim_{n\to\infty}\mathds{E}\|{\widetilde{\boldsymbol{\scriptstyle\mathcal{W}}}}_{n}\|^{2}\leq\|\mathcal{V}\|^{2}\mathds{E}\|{\mathcal{V}^{\sf T}\widetilde{\boldsymbol{\scriptstyle\mathcal{W}}}}_{n}\|^{2}=\|\mathcal{V}\|^{2}\mathds{E}(\|\bar{\boldsymbol{w}}_{n}\|^{2}+\|\check{\boldsymbol{w}}_{n}\|^{2})\leq O(\frac{\mu}{B})+O(\mu^{2})=O(\mu^{\gamma})$ (589)
@@
where
$\lim_{n\to\infty}(1-\mu\nu+O(\mu^{2}))^{n+1}=0$ (591)
for sufficiently small $\mu$.
H-B Proof for $\mathds{E}\|{\widetilde{\boldsymbol{\scriptstyle\mathcal{W}}}}_{n}\|^{4}$.
The derivation process is similar to Appendix C. The only difference is in (C) for the strongly-convex case, if we choose $t=1-\mu\nu$, it holds that
@@
$\lim_{n\to\infty}\mathds{E}\|{\widetilde{\boldsymbol{\scriptstyle\mathcal{W}}}}_{n}\|^{4}\leq O(\mu^{2\gamma})$ (620)
Similarly, for the centralized method, we substitute (590) into (H-B) and iterate it, and obtain
@@
from which for $n\to\infty$, the centralized method satisfies
$\lim_{n\to\infty}\mathds{E}\|{\widetilde{\boldsymbol{\scriptstyle\mathcal{W}}}}_{n}\|^{4}\leq O(\frac{\mu^{2}}{B^{2}})=O(\mu^{2(1+\eta)})$ (622)
H-C Proof for $|\mathds{E}\|\widetilde{\boldsymbol{\scriptstyle\mathcal{W}}}_{n}^{\prime}\|^{2}-\mathds{E}\|\widetilde{\boldsymbol{\scriptstyle\mathcal{W}}}_{n}\|^{2}|$ and $\mathds{E}\|{\widetilde{\boldsymbol{\scriptstyle\mathcal{W}}}}_{n}^{\prime}\|^{2}$.
We finally analyze the bounds of $|\mathds{E}\|\widetilde{\boldsymbol{\scriptstyle\mathcal{W}}}_{n}^{\prime}\|^{2}-\mathds{E}\|\widetilde{\boldsymbol{\scriptstyle\mathcal{W}}}_{n}\|^{2}|$ and $\mathds{E}\|{\widetilde{\boldsymbol{\scriptstyle\mathcal{W}}}}_{n}^{\prime}\|^{2}$. The proofs are similar to Appendix D. The only difference is that we now consider the long-term properties of $\mathds{E}\|\widetilde{\boldsymbol{\scriptstyle\mathcal{W}}}_{n}^{\prime}\|^{2}-\mathds{E}\|\widetilde{\boldsymbol{\scriptstyle\mathcal{W}}}_{n}\|^{2}|$ when $n\to\infty$. Similar to (312) and (317), after enough iterations when $n$ is sufficiently large, and with (322) and (620), we have
$\lim\limits_{n\to\infty}\mathds{E}\left[\begin{array}[]{c}\|\bar{\boldsymbol{z}}_{n}\|^{2}\\ \|\check{\boldsymbol{z}}_{n}\|^{2}\end{array}\right]\leq(I-\Gamma^{\prime})^{-1}\left[\begin{array}[]{c}O(\mu^{2\gamma+1})\\ O(\mu^{2\gamma+2})\end{array}\right]\leq\left[\begin{array}[]{cc}O(\frac{1}{\mu})&O(1)\\ O(\mu)&O(1)\end{array}\right]\left[\begin{array}[]{c}O(\mu^{2\gamma+1})\\ O(\mu^{2\gamma+2})\end{array}\right]=O(\mu^{2\gamma})$ (631)